# Semantic-Aware Spectrum Sharing in Internet of Vehicles Based on Deep Reinforcement Learning

### v3

@@ V Conclusion
In this paper, we introduced semantic information into high-speed mobile IoV environments, proposing the SSS scheme that utilizes the SAC DRL algorithm to maximize HSSE and the success probability of transmitting semantic data payloads. Simulation results illustrated the superior performance of the SSS algorithm to baseline spectrum sharing methods. The main conclusions are as follows:
  • HSSE initially increases and stabilizes with the growth of V2I link transmission power or the number of vehicles but decreases with increasing semantic demand sizes.
  • The success probability of transmitting effective semantic data payloads decreases with the increase in semantic demand.
  • When ${u_{q}}$ is less than 17 bits/word, indicating the need to transmit an average word length exceeding 17 bits in traditional communication systems, our semantic information transmission outperforms traditional bit-based systems. This highlights the importance of source coding schemes and the flexibility and efficiency of our proposed algorithm.

### v4

@@ V Conclusion
In this paper, we introduced semantic information into high-speed mobile IoV environments, proposing the SSS scheme that utilizes the SAC DRL algorithm to maximize HSSE and SRS. Simulation results illustrated the superior performance of the SSS algorithm to baseline spectrum sharing methods. The main conclusions are as follows:
  • HSSE increases and stabilizes with higher V2I link transmission power, increases with a higher number of vehicles, but decreases with larger semantic demand sizes.
  • The success probability of transmitting effective semantic data payloads decreases as semantic demand increases.
  • When ${u_{q}}$ is more than 17 bits/word, indicating the need to transmit an average word length exceeding 17 bits in traditional communication systems, our semantic information transmission outperforms traditional bit-based systems. This highlights the efficiency and flexibility of our proposed algorithm.
In future work, we will further explore how to deploy and test our algorithm in real-world scenarios.